How do I write and solve an equation for: the number of elephants at the circus was 2/5 of the number of horses. If there were 16 elephants, how many horses were there?
Is the equation 2/5x = 16? How do I solve it?
Thank you

Found 2 solutions by mathtutor19, solver91311:
Answer by mathtutor19(8)

That is the right equation.
2/5X=16
Multiply both sides by the reciprocal of 2/5, which is 5/2.
(5/2)•(2/5)X=16•(5/2)
The 5/2 and the 2/5 cancel out so you are left with:
X=16•(5/2)
Then you multiply 16 and 5/2 to get 40 so the answer is:
X=40
or there are 40 horses at the circus

Answer by solver91311(16868)



"Is the equation 2/5x = 16?"

Maybe. It depends on whether 2/5x = 16 means (the right way) or (the wrong way).

(2/5)x = 16 translates to

2/(5x) = 16 translates to

But

2/5x = 16 is ambiguous.

Presuming you meant to write (2/5)x = 16, just multiply both sides by to solve.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism